# Entropy Sentinel: Continuous LLM Accuracy Monitoring from Decoding Entropy Traces in STEM

### v2

@@ 5 Experimental Setup
6 Results
Research Questions. We organize results around four questions: (RQ1) does the entropy-profile signal support cross-domain accuracy estimation? (RQ2) how does it compare with simple baselines? (RQ3) how sensitive is performance to which benchmarks are used for supervision? (RQ4) how sensitive is performance to estimator design choices (classifier family, calibration, balancing)?
RQ1: Accuracy estimation under deployment-plausible defaults.
We first report two training configurations chosen a priori for plausibility rather than tuned for best score. Results here use a random forest correctness estimator trained with class balancing and isotonic calibration on Extremes (GSM8K + OlympiadBench), spanning elementary-to-competition difficulty, and Intermediate (MATH + SciBench) Table 2 and Figure 3 report cross-domain accuracy estimation quality.
First, difficulty-spanning supervision is consistently stronger: under Extremes, seven of nine models achieve $\rho\geq 0.90$ with low error (AEE 0.030.030.120.12), while Intermediate yields systematically higher error and weaker ordering (AEE 0.060.060.170.17; $\rho$ drops for all models). Second, the signal is model-dependent: for Phi-3.5-Mini we observe near-perfect ordering (AEE $0.03$, $\rho{=}1.00$), whereas Qwen3-8B exhibits weaker agreement in both settings (AEE 0.120.120.170.17, $\rho\approx 0.75$).
Overall, both choices generalize out of domain for most LLMs (Table 2). However, Extremes is consistently stronger, achieving lower AEE and higher rank agreement $\rho$ across models, while Intermediate incurs a systematic degradation. Additional examples under alternative training groups and estimator variants are reported in Appendix D.
RQ2: Baseline UQ metrics for performance monitoring.
We compare our entropy-based accuracy estimator against nine standard white-box uncertainty metrics in Sec. 3, defined in Appendix A. Because these metrics are not trained to predict accuracy (they return an uncertainty score rather than a calibrated correctness probability), we evaluate them in the most comparable way for monitoring: for each held-out domain, we aggregate each metric over instances to obtain a domain-level score, and report Spearman correlation $\rho$ between that score and the domain’s true accuracy. Table 4 summarizes results across models. In terms of raw correlation values, the best white-box metrics are in the same ballpark as our estimator—often within a few $\rho$ points—so they can track domain-to-domain fluctuations reasonably well. That said, our method is consistently as good as, and in most settings slightly better than, the strongest uncertainty baselines in Table 4, and it additionally supports a priori monitoring with calibrated, domain-level accuracy estimates as reported in Table 2; the baselines do not directly yield comparable predictions there, since they produce only uncalibrated scores.
Model Ext. Int. SEA $\text{SE}_{\text{max}}$ $\text{SE}_{\text{mean}}$ $\text{NLL}_{\text{avg}}$ $\text{NLL}_{\text{max}}$ LNTP MTP PPL
phi-3.5-mini (3B) 1.00 0.95 0.95 0.94 0.90 0.90 0.88 0.90 0.89 0.90
qwen3 (4B) 0.95 0.94 0.96 0.98 0.95 0.95 0.96 0.95 0.98 0.95
qwen3 (8B) 0.76 0.75 0.93 0.79 0.75 0.75 0.89 0.75 0.95 0.75
ministral3 (3B) 0.96 0.79 0.96 0.95 0.92 0.68 0.61 0.66 0.55 0.68
ministral3 (8B) 0.96 0.89 0.98 0.96 0.96 0.66 0.47 0.62 0.73 0.73
gemma3 (4B) 0.94 0.92 0.89 0.89 0.87 0.87 0.89 0.87 0.98 0.87
gemma3 (12B) 0.92 0.85 0.90 0.88 0.81 0.84 0.93 0.84 0.95 0.83
llama-3.1 (8B) 0.95 0.92 0.95 0.95 0.55 0.65 0.90 0.65 1.00 0.65
gpt-oss (20B) 0.90 0.89 0.94 0.98 0.82 0.58 0.37 0.50 0.70 0.50
Table 4: Spearman $\rho$ between aggregated UQ baselines and true domain accuracy on held-out benchmarks.GSM8K+OlympiadBench; Int.: MATH+SciBench. SEASEA and $SE_{\text{max}}$ are consistently strong, often comparable to our defaults.
RQ3: Training composition dominates (and improves with $k$).
We next quantify how accuracy estimation depends on the composition of the supervision set. Table 3 aggregates results over all training groups of size $k\in\{1,2,3,4\}$ (and over classifier configurations), reporting the median AEE and its interquartile range (IQR). Increasing $k$ produces two consistent effects across all nine LLMs: typical error decreases monotonically with $k$, and robustness to benchmark choice improves sharply (IQR shrinks). At k=1k{=}1, median AEE is high (roughly 0.200.200.270.27) and highly variable across which single benchmark provides supervision; by k=4k{=}4, median AEE improves substantially (roughly 0.060.060.160.16) and variability compresses to IQR $\approx 0.03$$0.04$. Thus, benchmark choice is a first-order design decision at small $k$, but becomes less brittle as supervision spans more tasks.
Difficulty balance explains much of the composition effect. To make the composition dependence interpretable, we summarize each training group by its weighted average accuracy (the average $A(D)$ over $D\in G$, weighted by instance counts) and relate it to held-out estimation quality. Figure 4 (and Appendix F for all models) shows a U-shaped relationship: supervision sets that are too easy or too hard generalize worse, whereas groups with intermediate weighted accuracy (roughly 0.40.40.60.6) yield the lowest AEE. A interpretation is that intermediate-weighted groups tend to mix easy and hard benchmarks, exposing the estimator to both low-entropy success patterns and high-entropy failure patterns; in contrast, easy-only or hard-only groups underrepresent one side of this spectrum and miscalibrate when transferred to unseen difficulty regimes. Difficulty-diverse training sets outperform difficulty-homogeneous ones: all-easy groups underrepresent high-entropy failure patterns from hard domains, while all-hard groups underrepresent low-entropy success patterns. Mixed (easy+hard) groups expose the estimator to a broader range of entropy profiles, which aligns with better OOD generalization. We analyze the best/worst benchmark combinations and a leave-one-out sensitivity study in Appendix F.
RQ4: Estimator choices matter, but less than data composition.
We isolate estimator design effects by aggregating across training groups and models. Table 5 reports main effects on median AEE and Spearman correlation. Random forests perform best in this low-dimensional setting (11 entropy features); logistic regression is close, while MLPs are consistently worse, consistent with overfitting under limited supervision. Isotonic calibration provides a modest but consistent gain (0.12 vs. 0.14 median AEE), which matters because we aggregate probabilities. Class balancing has little net effect and can slightly hurt, suggesting failures are driven more by entropy-pattern shift than label imbalance. Across all configurations, Spearman remains stable ($\rho\approx 0.94$), indicating the ranking of benchmark difficulty is robust and that most variation comes from training-data composition rather than estimator choices.
Feature Selection. Across ablations the full 11D profile is a strong, stable default, but it is not universally best. In particular, some smaller combinations often match or improve AEE and frequently improve $\rho$, indicating substantial redundancy among the summary statistics. More in Appendix E.
@@ 6 Results
MLP .14.07 .94.06
Calib. Y / N .12.07 / .14.07 .94.07 / .94.06
Bal. Y / N .13.06 / .13.07 .94.06 / .94.06
Table 5: Ablation main effects: median AEE and Spearman $\rho$ (IQR as subscripts). RF=Random Forest, LR=Logistic Regression.
@@
The most important design choice, above predictor or feature selection is the diversity of the supervision set. Training groups that span difficulty (mixing easy and hard tasks) consistently outperform difficulty-homogeneous groups, which tend to underrepresent either high-entropy failure patterns (easy-only) or low-entropy success patterns (hard-only).
Figure 4: Training-group difficulty vs. estimation quality for Phi-3.5-Mini: intermediate weighted accuracy (0.4–0.6) yields the lowest AEE; all-easy/all-hard groups perform worse.
@@ 7 Related Work
Supervised Approaches Using Internal Signals. Probe-based methods Belinkov (2022); Alain and Bengio (2017) train simple classifiers on model representations to predict properties of interest. For correctness prediction, prior work trains probes on LLM hidden states Azaria and Mitchell (2023); Chen et al. (2024); Zhang et al. (2025a), and ensembles combine multiple uncertainty signals Chen and Mueller (2023); Verga et al. (2024); Bouchard et al. (2025a). We likewise use a supervised probe, but target domain-level performance and rely only on final-layer entropy summaries, making the method more efficient and applicable to most open and closed-source models.
Entropy-Lens. Our focus on entropy signals is motivated by Entropy-Lens Ali et al. (2025), which uses Shannon entropy traces across transformer layers to classify model family, task type, and correctness. While effective, full residual-stream analysis is costly and yields high-dimensional features that may generalize poorly OOD. In contrast, we train on compact final-layer entropy summaries.
@@ 9 Limitations
Top-$k$ entropy is an approximation. We approximate Shannon entropy using only top-$k$ probabilities (top-20), omitting tail mass. This truncation can change the scale and shape of entropy traces, especially for high-entropy steps where probability mass is more diffuse. While this choice is motivated by API constraints, it may degrade performance compared to full-vocab entropy and can vary across tokenizers and model families.
@@
Model dependence and post-training effects. We observe that reliability varies across LLMs and can differ even within a family (e.g., size variants). One plausible factor is post-training (instruction tuning, RLHF/RLAIF, safety finetuning), which can decouple confidence signals from correctness. Consequently, our findings do not guarantee that entropy profiles will yield well-calibrated accuracy estimates for a new target model without empirical validation.
@@ Appendix A Baseline White Box UQ Metrics
We compare against nine standard uncertainty quantification baseline metrics derived exclusively from output LLM logprobs generated in regular inference. Most can be found in the uqlm library Bouchard et al. (2025b). These metrics are our most direct comparison in terms of efficient signals usable for potential continuous performance evaluation.
@@
Manakul et al. (2023) use $\text{NLL}(y_{i})=-\log P(y_{i}\mid x,y_{<i})$, aggregated as: $\text{NLL}_{\text{avg}}(y)=\frac{1}{L}\sum_{i=1}^{L}\text{NLL}(y_{i})$, $\text{NLL}_{\text{max}}(y)=\max_{i}\text{NLL}(y_{i})$, and $\text{NLL}_{\text{sum}}(y)=\sum_{i=1}^{L}\text{NLL}(y_{i})$. Note that the average and maximum metrics are covered in the compact vector presented in section 4.
@@ C.3 Classifier Hyperparameter Grid
All hyperparameters are selected via a 5 fold cross-validated grid search on the training group, optimizing ROC-AUC. For logistic regression ($\ell_{1}$ penalty, liblinear solver), we search $C\in\{0.5,2.0,10.0\}$. For random forests (100 estimators), we search max depth $\in\{3,5,10\}$ and min samples split $\in\{2,5,10\}$. For MLPs (ReLU activations, early stopping, $\alpha=0.001$), we search hidden layer sizes $\in\{(5),(8),(10),(15),(20),(8,4),(10,5),(15,8)\}$.
@@ Results.
Appendix E Reduced-Feature set Classifier
A striking finding is how little performance varies across feature sets. Differences in AEE between the full 11D profile and reduced alternatives typically fall within 0.01–0.03, and in many cases simpler feature sets match or outperform the full profile. For instance, SEA alone achieves lower AEE than the full profile for llama3-8B (0.05 vs. 0.07), ministral3-8B (0.06 vs. 0.07), and oss-20B (0.10 vs. 0.15) under Extremes supervision. Domain ranking quality ($\rho$) proves even more robust—for qwen3-8B, the full profile yields $\rho=0.76$ while SEA alone reaches $\rho=0.93$, suggesting additional features may introduce noise for models with weaker entropy–correctness coupling regarding difficulty ranking.
These results likely reflect high correlation among summary statistics—max, mean, quantiles, and accumulative metrics all derive from the same entropy trajectory, limiting the marginal information from additional features. Comparable results across feature sets supports our observation that training composition is the main driver of method performance.
Model Full 11D Max SEA Top2 S. Baselines
(AEE / $\rho$) (Default) (only) (only) (Max+SEA) (3-feat)
extremes (Random Forest; GSM + OlympiadBench)
Qwen3-8B 0.12 / 0.76 0.25 / 0.79 0.18 / 0.93 0.18 / 0.89 0.17 / 0.93
Phi3-3B 0.03 / 1.00 0.05 / 0.94 0.08 / 0.89 0.07 / 0.89 0.06 / 0.96
Qwen3-4B 0.08 / 0.95 0.08 / 0.98 0.09 / 0.98 0.08 / 0.98 0.09 / 0.96
Ministral3-3B 0.06 / 0.96 0.07 / 0.94 0.06 / 0.98 0.06 / 0.98 0.05 / 0.96
Ministral3-8B 0.07 / 0.96 0.08 / 0.95 0.06 / 0.98 0.06 / 0.98 0.07 / 0.98
Llama3-8B 0.07 / 0.95 0.07 / 0.95 0.05 / 0.94 0.05 / 0.94 0.05 / 0.94
Gemma3-4B 0.09 / 0.94 0.13 / 0.89 0.12 / 0.88 0.11 / 0.88 0.11 / 0.89
Gemma3-12B 0.08 / 0.92 0.12 / 0.88 0.09 / 0.90 0.09 / 0.88 0.09 / 0.88
OSS-20B 0.15 / 0.90 0.13 / 0.95 0.10 / 0.94 0.10 / 0.94 0.10 / 0.94
intermediate (Random Forest; MATH + SciBench)
Qwen3-8B 0.17 / 0.75 0.25 / 0.72 0.20 / 0.93 0.20 / 0.89 0.20 / 0.88
Phi3-3B 0.06 / 0.95 0.08 / 0.94 0.10 / 0.89 0.08 / 0.92 0.07 / 0.95
Qwen3-4B 0.11 / 0.94 0.09 / 0.98 0.10 / 0.96 0.10 / 0.98 0.10 / 0.96
Ministral3-3B 0.11 / 0.79 0.08 / 0.95 0.10 / 0.96 0.08 / 0.96 0.08 / 0.98
Ministral3-8B 0.12 / 0.89 0.09 / 0.96 0.11 / 0.98 0.09 / 0.96 0.09 / 0.96
Llama3-8B 0.11 / 0.92 0.11 / 0.94 0.07 / 0.95 0.07 / 0.94 0.07 / 0.94
Gemma3-4B 0.14 / 0.92 0.14 / 0.89 0.14 / 0.89 0.13 / 0.89 0.14 / 0.87
Gemma3-12B 0.12 / 0.85 0.13 / 0.88 0.11 / 0.90 0.11 / 0.88 0.12 / 0.88
OSS-20B 0.16 / 0.89 0.14 / 0.94 0.08 / 0.96 0.08 / 0.96 0.10 / 0.95
cross-domain (Log. Reg; MatSci + GSM-Symbolic)
Qwen3-8B 0.16 / 0.78 0.29 / 0.75 0.22 / 0.89 0.22 / 0.89 0.24 / 0.77
Phi3-3B 0.06 / 0.95 0.09 / 0.94 0.06 / 0.90 0.06 / 0.94 0.06 / 0.93
Qwen3-4B 0.08 / 0.96 0.11 / 0.98 0.08 / 0.98 0.09 / 0.98 0.08 / 0.98
Ministral3-3B 0.09 / 0.98 0.11 / 0.99 0.10 / 0.96 0.10 / 0.96 0.10 / 0.98
Ministral3-8B 0.08 / 0.95 0.12 / 0.96 0.11 / 0.98 0.11 / 0.96 0.11 / 0.95
Llama3-8B 0.09 / 0.95 0.12 / 0.95 0.06 / 0.95 0.06 / 0.95 0.05 / 0.96
Gemma3-4B 0.13 / 0.92 0.17 / 0.89 0.15 / 0.88 0.14 / 0.88 0.14 / 0.87
Gemma3-12B 0.10 / 0.88 0.16 / 0.88 0.10 / 0.90 0.10 / 0.90 0.10 / 0.90
OSS-20B 0.16 / 0.94 0.21 / 0.99 0.13 / 0.94 0.13 / 0.94 0.12 / 0.94
sci-math (MLP; GPQA + MatSci + GSM)
Qwen3-8B 0.19 / 0.78 0.23 / 0.73 0.18 / 0.95 0.18 / 0.89 0.19 / 0.88
Phi3-3B 0.07 / 0.95 0.08 / 0.94 0.06 / 0.90 0.06 / 0.90 0.07 / 0.93
Qwen3-4B 0.10 / 0.98 0.09 / 0.98 0.09 / 0.98 0.09 / 0.98 0.10 / 0.98
Ministral3-3B 0.07 / 0.95 0.06 / 0.98 0.08 / 0.96 0.06 / 0.96 0.06 / 0.96
Ministral3-8B 0.09 / 0.95 0.09 / 0.96 0.11 / 0.98 0.09 / 0.98 0.09 / 0.96
Llama3-8B 0.08 / 0.94 0.08 / 0.95 0.05 / 0.95 0.05 / 0.95 0.05 / 0.95
Gemma3-4B 0.13 / 0.92 0.12 / 0.89 0.11 / 0.89 0.11 / 0.89 0.11 / 0.89
Gemma3-12B 0.13 / 0.88 0.13 / 0.88 0.11 / 0.90 0.12 / 0.88 0.12 / 0.88
OSS-20B 0.15 / 0.93 0.16 / 0.96 0.13 / 0.98 0.13 / 0.98 0.13 / 0.98
Table 8: Cross-domain accuracy estimation across four supervision configurations and five feature sets. AEE: Median Absolute Estimation Error; $\rho$: Median Spearman correlation. Bold indicates best AEE (lowest) and best $\rho$ (highest) per row.
@@ Appendix F Additional Training Sensibility Results
U-Shaped Difficulty Relationship Across All Models
Figure 5 extends the analysis from Section 6 by plotting training group difficulty against estimation quality aggregated across all nine LLMs. The U-shaped pattern observed for Phi-3.5-Mini holds consistently: training groups with intermediate weighted accuracy (0.4–0.6) achieve the lowest median AEE, while groups at difficulty extremes yield degraded estimation. This reinforces that difficulty diversity in training data is more important than domain diversity for robust accuracy estimation.
@@
Table 9 identifies the highest- and lowest-performing benchmark combinations at each $k$, along with their weighted average group accuracy. The results directly corroborate the U-shaped relationship from Figures 4 and 5: best-performing groups fall within the intermediate accuracy regime (0.410.410.640.64), while worst-performing groups cluster at the low-accuracy extreme (0.110.110.220.22).

### v3

@@ 5 Results
RQ4: Sensitivity to estimator design.
We isolate estimator design effects by aggregating across all training groups and models. Table 5 reports main effects on median AEE and Spearman correlation across classifier architecture, calibration, balancing, and feature subsets.
Differences across all design axes are small: classifier choice and feature subset each shift median AEE by at most 0.030.030.040.04 points, calibration and balancing have near-zero effect, and Spearman remains stable at $\rho\approx 0.94$ throughout. Among the modest differences that do emerge, random forests with the 10D entropy-distribution statistics perform best—consistent with the plausible defaults chosen in RQ1—while MLPs show slight degradation, likely due to overfitting under limited supervision. The 17D profile performs slightly worse than 10D, suggesting the additional white-box UQ features introduce noise rather than complementary signal in this setting. Reducing to 3D or even 1D ($\text{SE}_{\text{sum}}$ alone) remains competitive, aligning with the RQ2 finding that accumulation and extrema metrics already capture most of the useful signal.
Overall, estimator design is a second-order concern: the best and worst classifier–feature combinations differ by only $0.04$ in median AEE, whereas training composition (RQ3) shifts median AEE by up to 0.100.100.150.15. This confirms the pattern suggested in RQ2, where simple Platt-scaled baselines under the same Extremes training group approached the multi-feature estimator—the choice of supervision data dominates both feature dimensionality and classifier complexity.
@@
MLP .15.08 .94.06
Calibration Y / N .13.07 / .13.08 .94.06 / .94.06
Balancing Y / N .13.07 / .13.08 .94.06 / .94.06
Features 17D .14.08 .94.06
10D .11.06 .95.05
3D .13.07 .94.07
1D .13.07 .94.06
Table 5: Ablation main effects: median AEE and Spearman $\rho$ (IQR as subscripts), aggregated across all training groups and models. RF=Random Forest, LR=Logistic Regression, MLP=Multi Layer Perceptron.
@@
The most important design choice, above predictor or feature selection, is the diversity of the supervision set. Training groups that span difficulty (mixing easy and hard tasks) consistently outperform difficulty-homogeneous groups, which tend to underrepresent either high-entropy failure patterns (easy-only) or low-entropy success patterns (hard-only).
@@ 6 Related Work
Supervised Approaches Using Internal Signals. Probe-based methods Belinkov (2022); Alain and Bengio (2017) train simple classifiers on model representations to predict properties of interest. For correctness prediction, prior work trains probes on LLM hidden states Azaria and Mitchell (2023); Chen et al. (2024); Zhang et al. (2025), and ensembles combine multiple uncertainty signals Chen and Mueller (2023); Verga et al. (2024); Bouchard and Chauhan (2025a). We likewise use a supervised probe, but target domain-level performance and rely only on final-layer entropy summaries, making the method more efficient and applicable to most open and closed-source models.
Entropy-Lens. Our focus on entropy signals is motivated primarily by Entropy-Lens Ali et al. (2025), which uses Shannon entropy traces across transformer layers to classify model family, task type, and correctness. While effective, full residual-stream analysis is costly and yields high-dimensional features that may generalize poorly OOD. In contrast, we train on compact final-layer entropy summaries.
@@ Appendix A Baseline White Box UQ Metrics
We compare against nine standard uncertainty quantification baseline metrics derived exclusively from output LLM logprobs generated in regular inference. Most can be found in the uqlm library Bouchard and Chauhan (2025b). These metrics are our most direct comparison in terms of efficient signals usable for potential continuous performance evaluation.
@@
Manakul et al. (2023) use $\text{NLL}(y_{i})=-\log P(y_{i}\mid x,y_{<i})$, aggregated as: $\text{NLL}_{\text{avg}}(y)=\frac{1}{L}\sum_{i=1}^{L}\text{NLL}(y_{i})$, $\text{NLL}_{\text{max}}(y)=\max_{i}\text{NLL}(y_{i})$, and $\text{NLL}_{\text{sum}}(y)=\sum_{i=1}^{L}\text{NLL}(y_{i})$.
@@ C.3 Classifier Hyperparameter Grid
All hyperparameters are selected via a 5 fold cross-validated grid search on the training group, optimizing ROC-AUC. For the logistic regression, we don’t train for any hyperparameter. For random forests (100 estimators), we search max depth $\in\{3,5,10\}$ and min samples split $\in\{2,5,10\}$. For MLPs (ReLU activations, early stopping, $\alpha=0.001$), we search hidden layer sizes $\in\{(5),(8),(10),(15),(20),(8,4),(10,5),(15,8)\}$.
@@ Results.
Appendix E Full Baseline Results
Full baseline results can be found in Table 8.
Model Default $\text{SE}_{\text{sum}}$ $\text{SE}_{\text{max}}$ $\text{SE}_{\text{mean}}$ $\text{NLL}_{\text{avg}}$ $\text{NLL}_{\text{max}}$ $\text{NLL}_{\text{sum}}$ LNTP MTP PPL
phi-3.5-mini (3B) $\rho$ 1.00 0.90 0.94 0.93 0.93 0.89 0.98 0.86 0.89 0.93
AEE 0.03 0.08 0.06 0.12 0.13 0.12 0.08 0.13 0.12 0.13
qwen3 (4B) $\rho$ 0.95 0.98 0.96 0.96 0.96 0.96 0.98 0.95 0.98 0.96
AEE 0.08 0.10 0.07 0.13 0.13 0.11 0.10 0.11 0.12 0.13
qwen3 (8B) $\rho$ 0.76 0.92 0.75 0.75 0.77 0.90 0.92 0.82 0.90 0.77
AEE 0.12 0.18 0.25 0.22 0.22 0.27 0.19 0.25 0.27 0.22
ministral3 (3B) $\rho$ 0.96 0.98 0.95 0.96 0.73 0.56 0.96 0.14 0.61 0.73
AEE 0.06 0.07 0.07 0.09 0.18 0.29 0.07 0.30 0.29 0.19
ministral3 (8B) $\rho$ 0.96 0.98 0.95 0.96 0.73 0.47 0.98 0.61 0.47 0.77
AEE 0.07 0.08 0.08 0.10 0.19 0.27 0.08 0.20 0.27 0.19
gemma3 (4B) $\rho$ 0.94 0.88 0.89 0.84 0.87 0.92 0.88 0.38 0.92 0.87
AEE 0.09 0.13 0.14 0.20 0.21 0.22 0.12 0.29 0.24 0.21
gemma3 (12B) $\rho$ 0.92 0.90 0.88 0.84 0.84 0.93 0.90 0.88 0.93 0.84
AEE 0.08 0.11 0.13 0.20 0.21 0.17 0.11 0.18 0.18 0.21
llama-3.1 (8B) $\rho$ 0.95 0.95 0.95 0.61 0.61 0.99 0.96 0.36 0.99 0.61
AEE 0.07 0.06 0.08 0.28 0.28 0.13 0.06 0.29 0.14 0.28
gpt-oss (20B) $\rho$ 0.90 0.94 0.95 0.75 0.47 0.42 0.94 0.60 0.42 0.50
AEE 0.15 0.10 0.13 0.17 0.22 0.23 0.10 0.23 0.23 0.21
Table 8: Spearman $\rho$ (higher is better) and AEE (lower is better) between UQ baselines and true domain accuracy on held-out benchmarks.
@@ Appendix F Additional Training Sensibility Results
Table 9 identifies the highest- and lowest-performing benchmark combinations at each $k$, along with their weighted average group accuracy. The results directly corroborate the U-shaped relationship from Figure 4: best-performing groups fall within the intermediate accuracy regime (0.40.40.70.7).2